# Automated Test-Case Generation for REST APIs Using Model Inference Search Heuristic

### v1

@@ I Introduction
To address this limitation, we introduce the Model Inference Search Heuristic (MISH), a novel grey-box approach to guide the automatic test case generation process. Unlike white-box techniques that rely on code-level metrics such as branch distance, MISH leverages log statements printed by the SUT during test execution, which are inserted into the code by developers. MISH continuously collects log statements output by the SUT during test execution and uses them to learn a finite state automaton (or simply a state machine) in real time. Each test case is converted into a sequence of events derived from these logs, and the inferred state machine serves as a representation of the system’s behavior. We learn a state machine using the online learning heuristic proposed by Baumgartner and Verwer for FlexFringe [13], a framework for learning state machines. Our approach uses a single objective to define the fitness of a test case, the path traversed within the state machine. To the best of our knowledge, we are the first to apply real-time state machine learning to the task of automated test-case generation for microservices’ REST APIs.
As state machines capture sequential behaviors exhibited by an SUT, we can leverage this information to capture the required sequential pattern of REST API calls and create the corresponding sequential behavior within the SUT. This way, we can preserve sequential patterns that can be used to generate test cases for complex interconnected behaviors between microservices. We believe these sequential patterns will lead to better code coverage and fault detection.
@@ II-B Automated Testing of REST APIs
RESTTESTGEN [48], RESTest [36], and RapiTest [22] are methods developed to automatically test REST APIs that use black-box heuristics. EvoMaster also provides the option to generate test cases following in black-box mode [9]. These typically use the API specifications of the SUT to generate test cases. For the white-box approach, many state-of-the-art techniques are built as search algorithms within EvoMaster. The Many Independent Objective (MIO) algorithm is a state-of-the-art technique built within EvoMaster [6]. This algorithm treats each test target as an independent objective and tries to find a test case that covers the objective. The aim is to find a test suite that covers as many (independent) objectives as possible. MIO keeps track of a population of test cases for each objective and creates the final test suite by taking the union of the best test cases from all populations. The Many Objective Sorting Algorithm (MOSA) is another state-of-the-art technique built within EvoMaster. Proposed by Panichella et al. [38], this algorithm treats test case generation as a multi-objective optimization problem. Instead of optimizing each objective separately as with the MIO algorithm, all objectives are optimized at the same time in MOSA. Unlike MIO, MOSA keeps track of a single population of test cases. The final result is a minimal test suite containing the test cases covering as many objectives as possible.
@@ II-C State Machine & Testing REST APIs
Several works have shown the effectiveness of state machines in testing and verification of the behavior of different systems [54, 21, 23, 33, 37, 27]. However, the focus of these works was not on automated test-case generation for REST-APIs. Cao et al. have applied state machines on REST-APIs [17, 18] for different purposes. However, the focus of these two works was also not on automated test-case generation. Several works have specifically used state machines to generate test cases automatically for REST APIs. These approaches typically extract a state machine from the UML diagrams generated for the microservices [39, 53, 3]. These approaches assume that such diagrams are available for the SUT and are up to date with the latest version of the SUT.
@@ III-C Defining Fitness using State Machine
Once the state machine has been updated, MISH forwards the traces (generated from the test cases) to the model inference framework and requests it to compute their fitness (lines 7 in Algorithm 1). The fitness value computed for a particular trace is assigned as the fitness value for the corresponding test case. Fitness is determined based on the path (the sequence of states) traversed by the trace within the state machine. The primary objective of MISH is to generate system-level test cases that trigger deep and interconnected interactions between microservices. Each state in the state machine represents a unique state of the SUT, and its visit frequency is computed from the traces used to update the model. Test cases that lead the system into less frequently visited states (or traverse paths containing more of such states) are more likely to reveal complex and interconnected behaviors, thus contributing to higher system-level coverage. Consequently, MISH assigns higher fitness values to test cases that visit (or traverse paths through) infrequent states.
@@
Weighted State Visit Frequencies (WS): this function first collect the visit frequency of the states visited by an arbitrary trace $t$.These frequencies are sorted in ascending order. A weighted sum is then calculated by multiplying each state’s visit frequency by its index in the sorted list (starting from 1). The fitness value of the trace is defined as the inverse of this weighted sum. Using the same trace $t$ from Figure1, the sorted state visit frequencies are 6, 15, 15, and 15. The weighted sum of these frequencies is then $(6\times 1)+(15\times 2)+(15\times 3)+(15\times 4)=141$. The fitness of $t$ is then $\frac{1}{141}{\approx}0.007$.
@@ III-D Evolving and Selecting Test Cases
In each iteration, MISH generates an offspring population based on the current population of test cases (see line 9 in Algorithm 1). The offspring population has the same size as the current population. Specifically, MISH uses tournament selection to select candidate test cases for the offspring population. MISH then creates a copy of the selected candidate and introduces a mutation in the copied test case. The copied test case is then added to the offspring population. MISH selects 90% of the offspring through tournament selection and the remainder of the offspring population is filled by sampling test cases to create more diverse individuals for better exploration. We employ the same sampling strategy as MOSA. MISH then computes the fitness of the offspring population following the procedure described previously. Finally, MISH selects the new population (see line 16 of in Algorithm 1) of test cases for the next iteration by taking the top $S$ test cases from the union of the current population and the offspring population, where $S$ denotes the population size.
@@ IV-B Selection of Applications
To evaluate MISH’s effectiveness in generating system-level test cases, we benchmark MISH on six real-world microservice applications from the EMB dataset. This dataset is created by the same authors as EvoMaster and was specifically designed to benchmark test case generation approaches of EvoMaster [12]. The version number of the dataset is synced with the version number of EvoMaster. As MISH is developed within EvoMaster, we have selected this dataset for our evaluation. Specifically, we used version 2.0 of the dataset as this was the latest stable release at the time of our implementation. While version 3.2 of EvoMaster was released at the time of writing this work, we found it does not report a specific category of faults (Potential Faults) that we use to evaluate MISH’s effectiveness (see Subsection IV-E in the description of the metrics). Additionally, we encountered errors when attempting to write test cases to file. These two points led us to use an earlier version of the dataset for evaluation instead of the latest version. We selected only the applications in which REST APIs are used as MISH is designed for testing REST APIs. The six real-world applications that we have selected are:
@@ IV-D Execution Procedure of Each Algorithm
To answer our research questions, we ran each algorithm (MOSA and MISH) 20 times on each selected microservice application to mitigate the effects of stochastic operations (e.g., mutation) on performance results. Each run is executed independently, i.e., the application and the EA are reset after each run to make sure that they are both started in a clean state for collecting performance statistics. Since MISH implements two distinct fitness functions, each version of MISH was also run 20 times per application. In total, we executed 360 runs, which would take approximately 15 days to complete if run sequentially. To speed up the evaluation process, we created a Docker image that enables us to parallelize the runs. This was inspired by the experimental setup used by Arcuri et al. [10]. Each Docker container executes 20 runs of an algorithm on a particular application. We make our Docker image and source code publicly available[5] to promote the reproducibility of our study. Currently, our Docker image does not include the option to run OCVN as this application also starts up a Docker container. We chose not to run Docker containers within another Docker container due to potential performance and compatibility issues. For experiments involving OCVN, we executed all three algorithms directly on the host machine, ensuring a fair comparison by using the same hardware specifications for all runs.
@@ IV-E Comparison Metrics
To determine whether the results between the algorithms are significantly different, we performed unpaired Wilcoxon rank-sum tests with a threshold of 0.05 to compute the $\rho$-values. This is a non-parametric pairwise test for determining whether a pair of data distributions are significantly different from each other [19]. Since MOSA is our baseline method and we have two different versions of MISH, we performed a pairwise comparison between each version of MISH and MOSA. Furthermore, to determine the magnitude of the difference between the two versions MISH and MOSA, we computed the effect size of the result using the Vargha-Delaney statistic [47].
@@ V-A RQ1: MISH’s Performance In Generating Test Cases
Tables IIIIV, and V present the performance of each algorithm for every application according to the respective comparison metric. To address the outliers caused by random processes in EAs, we report the medians along with their corresponding interquartile ranges (IQR) instead of the mean.
From Table III, both versions of MISH achieve higher coverage than MOSA on Features-Service, ProxyPrint, and OCVN. On the remaining applications, MISH demonstrates coverage results that are competitive with MOSA. MISH achieved on average a minimal increase (0.68%) in coverage on Feature-Service when compared to MOSA. However, MISH achieved a much larger increase in coverage on ProxyPrint and OCVN (15.7% and 42.4%, respectively). Regarding fault detection capabilities (Tables IV and V), MISH again outperforms MOSA on ProxyPrint and OCVN, while detecting a comparable number of faults as MOSA on the other applications. Specifically, for the potential faults detected on ProxyPrint and OCVN, MISH detected (on average) 4.5% and 13.5% more faults than MOSA, respectively. For the potential 500 faults, MISH detected (on average) 6.3% more faults on ProxyPrint and 76.5% more faults on OCVN. Tables VIVII, and V present the results of the statistical tests conducted for each comparison metric. As expected, there is no statistical significance in MISH’s coverage performance on Feature-Service as it achieved only a minimal increase in coverage. However, MISH performed significantly better on ProxyPrint and OCVN across all three metrics, each associated with a large effect size. The results of our statistical tests confirm our findings presented in Tables IIIIV and V.
@@ V-B RQ2: Where does MISH exhibit good/poor performance?
To understand MISH superior performance on OCVN, we compared the faults identified by MISH-LM and MISH-WS with those identified by MOSA. Specifically, we analyzed the unique set of fault IDs detected across all runs this application. Figure 3 presents the Venn diagram on the faults detected by each algorithm in OCVN. As we can observe, each version of MISH has identified more unique faults than MOSA. Notably, MISH-LM identified more than twice as many faults as MOSA.
@@
An important observation from our evaluation is that MISH performs particularly well on applications with a larger number of API endpoints, such as ProxyPrint and OCVN. Additional investigation into the faults identified only by MISH on ProxyPrint revealed a pattern similar to OCVN: the first call logs in as an administrator, followed by a sequence of dependent API calls. We believe that MISH’s superior performance on these two applications is due to its ability to capture and exploit the dependencies between REST API calls, which MOSA (and its many-objective design) does not explicitly account for.
@@ V-C RQ3: MISH’s Coverage Over Time
The one exception is LanguageTool, where MISH does not outperform MOSA in discovering unseen test targets. On this application, MISH’s coverage performance reaches a plateau early in the run and remains stagnant for the rest of the execution. Two key factors contribute to MISH’s performance on LanguageTool. First, LanguageTool has the largest number of source files and lines of code among the applications, which increases the complexity of the system under test. Second, LanguageTool has only two API endpoints, limiting MISH’s ability to generate diverse traces. MISH relies on log statements produced by these endpoints to create traces and update the state machine. Furthermore, MISH can only modify the parameters and actions of the HTTP requests targeting these endpoints to discover new test targets. These constraints restrict MISH’s ability to explore deeper or more diverse behaviors, especially when the API endpoints operate independently and lack dependencies that could be leveraged for sequential patterns.
@@ VI Threat to Validity
Concerning threats to construct validity, we evaluated all evolutionary algorithms (EAs) using metrics widely employed in prior studies: the number of covered targets, the number of potential faults detected, and the number of potential 500 faults detected. For the stopping criterion, we utilized a commonly used approach based on a fixed search budget measured by runtime. Specifically, we set the search budget to 60 minutes per run, consistent with prior research on EvoMaster.
@@
To address threats to conclusion validity, we followed best practices from prior studies on EvoMaster. Each algorithm was run 20 times per application, using different random seeds to account for the stochastic nature of EAs. For statistical analysis, we employed the Wilcoxon rank-sum test to evaluate the significance of differences between algorithms and the Vargha-Delaney effect size measure to quantify the magnitude of these differences.

### v2

@@ I Introduction
To address this limitation, we introduce the Model Inference Search Heuristic (MISH), a novel grey-box approach to guide the automatic test case generation process. Unlike white-box techniques that rely on code-level metrics such as branch distance, MISH leverages log statements printed by the SUT during test execution, which developers insert into the code. MISH continuously collects log statements output by the SUT during test execution and uses them to learn a finite state automaton (or simply a state machine) in real-time. Each test case is converted into a sequence of events derived from these logs, and the inferred state machine serves as a representation of the system’s behavior. We learn a state machine using the online learning heuristic proposed by Baumgartner and Verwer for FlexFringe [12], a framework for learning state machines. Our approach uses a single objective to define the fitness of a test case, the path traversed within the state machine. To the best of our knowledge, we are the first to apply real-time state machine learning to the task of automated test-case generation for microservices’ REST APIs.
As state machines capture sequential behaviors exhibited by a SUT, we can leverage this information to capture the required sequential pattern of REST API calls and create the corresponding sequential behavior within the SUT. This way, we can preserve sequential patterns that can be used to generate test cases for complex interconnected behaviors between microservices. We believe these sequential patterns will lead to better code coverage and fault detection.
@@ II-B Automated Testing of REST APIs
RESTTESTGEN [50], RESTest [38], and RapiTest [24] are methods developed to automatically test REST APIs using black-box heuristics. EvoMaster also provides the option to generate test cases following in black-box mode [8]. These typically use the API specifications of the SUT to generate test cases. For the white-box approach, many state-of-the-art techniques are built as search algorithms within EvoMaster. The Many Independent Objective (MIO) algorithm is a state-of-the-art technique built within EvoMaster [5]. This algorithm treats each test target as an independent objective and tries to find a test case that covers the objective. The aim is to find a test suite that covers as many (independent) objectives as possible. MIO keeps track of a population of test cases for each objective and creates the final test suite by taking the union of the best test cases from all populations. The Many Objective Sorting Algorithm (MOSA) is another state-of-the-art technique built within EvoMaster. Proposed by Panichella et al. [40], this algorithm treats test case generation as a multi-objective optimization problem. Instead of optimizing each objective separately, as with the MIO algorithm, all objectives are optimized simultaneously in MOSA. Unlike MIO, MOSA keeps track of a single population of test cases. The final result is a minimal test suite containing the test cases covering as many objectives as possible.
@@ II-C State Machine & Testing REST APIs
Several works have shown the effectiveness of state machines in testing and verification of the behavior of different systems [56, 23, 25, 35, 39, 29]. However, the focus of these works was not on automated test-case generation for REST-APIs. Cao et al. have applied state machines on REST-APIs [17, 20, 16] for different purposes. However, the focus of these two works was also not on automated test-case generation. Several works have specifically used state machines to generate test cases automatically for REST APIs. These approaches typically extract a state machine from the UML diagrams generated for the microservices [41, 55, 3]. These approaches assume that such diagrams are available for the SUT and are up to date with the latest version of the SUT.
@@ III-C Defining Fitness using State Machine
Once the state machine has been updated, MISH forwards the traces (generated from the test cases) to the model inference framework and requests it to compute their fitness (line 7 in Algorithm 1). The fitness value computed for a particular trace is assigned as the fitness value for the corresponding test case. Fitness is determined based on the path (the sequence of states) traversed by the trace within the state machine. The primary objective of MISH is to generate system-level test cases that trigger deep and interconnected interactions between microservices. Each state in the state machine represents a unique state of the SUT, and its visit frequency is computed from the traces used to update the model. We assume that test cases that lead the system into less frequently visited states (or traverse paths containing more of such states) are more likely to reveal complex and interconnected behaviors, thus contributing to higher system-level coverage. Consequently, MISH assigns higher fitness values to test cases that visit (or traverse paths through) infrequent states.
@@
Weighted State Visit Frequencies (WS): this function first collects the visit frequency of the states visited by an arbitrary trace $t$. These frequencies are sorted in ascending order. A weighted sum is then calculated by multiplying each state’s visit frequency by its index in the sorted list (starting from 1). The fitness value of the trace is defined as the inverse of this weighted sum. Using the same trace $t$ from Figure 1, the sorted state visit frequencies are 6, 15, 15, and 15. The weighted sum of these frequencies is then $(6\times 1)+(15\times 2)+(15\times 3)+(15\times 4)=141$. The fitness of $t$ is then $\frac{1}{141}{\approx}0.007$.
@@ III-D Evolving and Selecting Test Cases
In each iteration, MISH generates an offspring population based on the current population of test cases (see line 9 in Algorithm 1). The offspring population has the same size as the current population. Specifically, MISH uses tournament selection to select candidate test cases for the offspring population. MISH then creates a copy of the selected candidate and introduces a mutation in the copied test case. The copied test case is then added to the offspring population. MISH selects 90% of the offspring through tournament selection, and the remainder of the offspring population is filled by sampling test cases to create more diverse individuals for better exploration. We employ the same sampling strategy as MOSA. MISH then computes the fitness of the offspring population following the procedure described previously. Finally, MISH selects the new population (see line 16 of in Algorithm 1) of test cases for the next iteration by taking the top $S$ test cases from the union of the current population and the offspring population, where $S$ denotes the population size.
@@ IV-B Selection of Applications
To evaluate MISH’s effectiveness in generating system-level test cases, we benchmark MISH on six real-world microservice applications from the EMB dataset. This dataset is created by the same authors as EvoMaster and was specifically designed to benchmark test case generation approaches of EvoMaster [11]. The version number of the dataset is synced with the version number of EvoMaster. As MISH is developed within EvoMaster, we have selected this dataset for our evaluation. Specifically, we used version 2.0 of the dataset as this was the latest stable release at the time of our implementation. While version 3.2 of EvoMaster was released at the time of writing this work, we found it does not report a specific category of faults (Potential Faults) that we use to evaluate MISH’s effectiveness (see Subsection IV-E in the description of the metrics). Additionally, we encountered errors when attempting to write test cases to file. These two points led us to use an earlier version of the dataset for evaluation instead of the latest version. We selected only the applications in which REST APIs are used, as MISH is designed to test REST APIs. The six real-world applications that we have selected are:
@@ IV-D Execution Procedure of Each Algorithm
To answer our research questions, we ran each algorithm (MOSA and MISH) 20 times on each selected microservice application to mitigate the effects of stochastic operations (e.g., mutation) on performance results. Each run is executed independently, i.e., the application and the EA are reset after each run to make sure that they are both started in a clean state for collecting performance statistics. Since MISH implements two distinct fitness functions, each version of MISH was also run 20 times per application. In total, we executed 360 runs, which would take approximately 15 days to complete if run sequentially. To speed up the evaluation process, we created a Docker image that enables us to parallelize the runs. This was inspired by the experimental setup used by Arcuri et al. [9]. Each Docker container executes 20 runs of an algorithm on a particular application. We make our Docker image and source code publicly available[18, 19] to promote the reproducibility of our study. Our Docker image does not include the option to run OCVN, as it spawns a Docker container. We chose not to run Docker containers within another Docker container due to potential performance and compatibility issues. For experiments involving OCVN, we executed all three algorithms directly on the host machine, ensuring a fair comparison by using the same hardware specifications for all runs.
@@ IV-E Comparison Metrics
To determine whether the results between the algorithms are significantly different, we performed unpaired Wilcoxon rank-sum tests with a threshold of 0.05 to compute the $\rho$-values. This is a non-parametric pairwise test for determining whether a pair of data distributions are significantly different from each other [21]. Since MOSA is our baseline method and we have two different versions of MISH, we performed a pairwise comparison between each version of MISH and MOSA. Furthermore, to determine the probability that a sample of results from one method will exceed those from the other method, we computed the effect size of the result using the Vargha-Delaney statistic [49]
@@ V-A RQ1: MISH’s Performance In Generating Test Cases
Tables IIIIV, and V present the performance of each algorithm for every application according to the respective comparison metric. The best-performing algorithm for each corresponding application is highlighted in bold. To address the outliers caused by random processes in EAs, we report the medians along with their corresponding interquartile ranges (IQR) instead of the mean.
From Table III, both versions of MISH achieve higher coverage than MOSA on Features-Service, ProxyPrint, and OCVN. On the remaining applications, MISH demonstrates coverage results that are competitive with MOSA. MISH achieved, on average, a minimal increase (0.68%) in coverage on Feature-Service when compared to MOSA. However, MISH achieved a much larger increase in coverage on ProxyPrint and OCVN (15.7% and 42.4%, respectively). Regarding fault detection capabilities (Tables IV and V), MISH again outperforms MOSA on ProxyPrint and OCVN, while detecting a comparable number of faults as MOSA on the other applications. Specifically, for the potential faults detected on ProxyPrint and OCVN, MISH detected (on average) 4.5% and 13.5% more faults than MOSA, respectively. For the potential 500 faults, MISH detected (on average) 6.3% more faults on ProxyPrint and 76.5% more faults on OCVN. Tables VIVII, and V present the results of the statistical tests conducted for each comparison metric. Significant $\rho$-values ($<0.05$) are highlighted in gray. As expected, there is no statistical significance in MISH’s coverage performance on Feature-Service as it achieved only a minimal increase in coverage. However, MISH performed significantly better on ProxyPrint and OCVN across all three metrics, each associated with a large effect size. The results of our statistical tests confirm our findings presented in Tables IIIIV and V.
@@ V-B RQ2: Where does MISH exhibit good/poor performance?
To understand MISH’s superior performance on OCVN, we compared the faults identified by MISH-LM and MISH-WS with those identified by MOSA. Specifically, we analyzed the unique set of fault IDs detected across all runs of this application. Figure 3 presents the Venn diagram on the faults detected by each algorithm in OCVN. As we can observe, each version of MISH has identified more unique faults than MOSA. Notably, MISH-LM identified more than twice as many faults as MOSA.
@@
An important observation from our evaluation is that MISH performs particularly well on applications with numerous API endpoints, such as ProxyPrint and OCVN. Additional investigation into the faults identified only by MISH on ProxyPrint revealed a pattern similar to OCVN: the first call logs in as an administrator, followed by a sequence of dependent API calls. We believe that MISH’s superior performance on these two applications is due to its ability to capture and exploit the dependencies between REST API calls, which MOSA (and its many-objective design) does not explicitly account for.
@@ V-C RQ3: MISH’s Coverage Over Time
The one exception is LanguageTool, where MISH does not outperform MOSA in discovering unseen test targets. On this application, MISH’s coverage performance reaches a plateau early in the run and remains stagnant for the rest of the execution. Two key factors contribute to MISH’s performance on LanguageTool. First, LanguageTool has the largest number of source files and lines of code among the applications, which increases the complexity of the system under test. Second, LanguageTool has only two API endpoints, limiting MISH’s ability to generate diverse traces. MISH relies on log statements produced by these endpoints to create traces and update the state machine. Furthermore, MISH can only adjust the HTTP request parameters and actions directed at these endpoints to identify new test targets. This limitation hinders MISH’s ability to uncover more varied behaviors, particularly when the API endpoints function independently and lack dependencies that might reveal sequential patterns.
@@ VI Threat to Validity
Concerning threats to construct validity, we evaluated all evolutionary algorithms (EAs) using metrics widely employed in prior studies: the number of covered targets, the number of potential faults detected, and the number of potential 500 faults detected. For the stopping criterion, we utilized a commonly used approach based on a fixed search budget measured by runtime. Specifically, we set the search budget to 60 minutes per run, consistent with prior EvoMaster research.
@@
To address threats to conclusion validity, we followed best practices from prior studies on EvoMaster. Each algorithm was run 20 times per application, using different random seeds to account for the stochastic nature of EAs. For statistical analysis, we employed the Wilcoxon rank-sum test to evaluate the significance of differences between algorithms and the Vargha-Delaney effect size measure to quantify their magnitude.